Fw: scheduling update and news - IR access.

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Wed Nov 07 2001 - 20:27:21 EST

  • Next message: Apache: "Shift report 20011108 16:00-24:00"

    Tomorrow,
     during the day time RHIC IR will be in restricted access until abpout 3-4 pm.We have little to do except for possibly
    a) checking the T5 status
    b) possible moving the MRS platform i.e. Tpm1,tpm2 and D5 back by ~ .5 m to check and possibly improve the signal/ background
    at angles like 35/40. The solid angles will be somewhat modified, but particle densities in tpm1 should be smaller
    - If it does not change performance radically we can go back shortly to default configuration.
    c) checking and fixing the MRS stuck at 90 degreees (happended last night)
    
    /FV
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    ----- Original Message ----- 
    From: Angelika Drees 
    Sent: Wednesday, November 07, 2001 5:21 PM
    Subject: scheduling update and news
    
    
    Hey everybody, 
    (don't get confused, it's wednesday!) Todays beam experiments are finished  pretty much on time and we are currently at injection and trying to refill. Unfortunately due to QLIs this morning and a booster main magnet problem earlier, some items - like the Vernier Scan :( - were sacrificed. However, at this point we have  problems with the Ramp Editor and the WFGs (unscheduled FEC reboots) preventing us from activating any changes. We are currently working on this. In order to save *some* beam time for the SEB experiments transition/octupole tuning for higher intensities (as Dejan suggested in the 9:00 meeting)  got postponed until tomorrow evening when we come back on from the access anyway. 
    The IRs will be in restricted access tomorrow. Please contact MCR for other sectors. Everybody has to be out by 4:00 pm, preferably earlier so we can start sweeping certain areas before hand.  Cabling for the AGS gammaT quad will be done as time allows. 
    
    One item slipped thru my mind when I presented the schedule yesterday: beta*=3m injection studies to prepare the  pp run.  I apologize. The schedule should read ~ 2-3  hours for this tuesday morning pending a RHIC store (monday is Veterans Day and some important "veterans" are not around!). 
    
    SEB finishes tonight :(. Please come to the next time meeting though to a) invite us to an end-of-run party b) present final notes/comments/results etc. SEB for E949 will resume later this (fiscal) year. 
    
    
    -- 
                                                    \|/
                     Angelika                      (.~.)
                                                  o0(_)0o
    -----------------------------------------------------------
    Angelika Drees (Collider-Accelerator Department, AP group)
    BNL, Building 911b-210
    work: (631) 344 2348
    FAX:  (631) 344 5954      
    home: (631) 474 3884
      
    



    This archive was generated by hypermail 2b30 : Wed Nov 07 2001 - 20:25:23 EST